On January 2, 2018, David Corporation purchased a patent for $700,000. The remaining legal life is 12 years, but the company estimated that the patent will be useful only for eight years. In January 2020, the company incurred legal fees of $105,000 in successfully defending a patent infringement suit. The successful defense did not change the company’s estimate of useful life.

Required:
Prepare journal entries related to the patent for 2018, 2019, and 2020. (If no entry is required for a transaction/event, select "No journal entry required" in the first account field.)
Answer #1
January 02, 2018 Patents 700000
      Cash 700000
December 31, 2018 Amortization expense 87500 =700000/8
      Patents 87500
December 31, 2019 Amortization expense 87500 =700000/8
      Patents 87500
January 2020 Patents 105000
      Cash 105000
December 31, 2020 Amortization expense 105000 =(700000-87500-87500+105000)/6
      Patents 105000